### v1

# Setting the Course, but Forgetting to Steer: Analyzing Compliance with GDPR’s Right of Access to Data by Instagram, TikTok, and YouTube

Abstract
The comprehensibility and reliability of data download packages (DDPs) provided under the General Data Protection Regulation’s (GDPR) right of access are vital for both individuals and researchers. These DDPs enable users to understand and control their personal data, yet issues like complexity and incomplete information often limit their utility. Also, despite their growing use in research to study emerging online phenomena, little attention has been given to systematically assessing the reliability and comprehensibility of DDPs.
To bridge this research gap, in this work, we perform a comparative analysis to assess the comprehensibility and reliability of DDPs provided by three major social media platforms, namely, TikTok, Instagram, and YouTube. By recruiting 400 participants across four countries, we assess the comprehensibility of DDPs across various requirements, including conciseness, transparency, intelligibility, and clear and plain language. Also, by leveraging automated bots and user-donated DDPs, we evaluate the reliability of DDPs across the three platforms. Among other things, we find notable differences across the three platforms in the data categories included in DDPs, inconsistencies in adherence to the GDPR requirements, and gaps in the reliability of the DDPs across platforms. Finally, using large language models, we demonstrate the feasibility of easily providing more comprehensible DDPs.
  *These authors contributed equally to this work.
1 Introduction
Ensuring compliance with the General Data Protection Regulation (GDPR) [12] is essential for platforms handling personal data, as it establishes legal requirements for data processing, security, and user rights. One of these rights is the right of access by data subjects, which enlists provisions whereby data subjects (i.e., end users) can request access to their data being processed by platforms (Article 15(3) GDPR). Platforms have to comply with these requests within a stipulated period of one month and provide the user with a copy of their data. Usually, platforms provide this data in the form of data download packages (DDPs) [34]. These DDPs often include information about the users’ activities on and off the platform, device and app information, inferred preferences, and other information depending on the business of the platform.
Implementation across platforms: In fact, the GDPR stipulates that information shall be provided to users in commonly used electronic forms (Art. 15(3)(2) GDPR). Such an implementation is crucial for enabling users to exercise their right to data portability (Article 20), which allows end users to transmit data from one platform to the other without hindrance [12]. However, the GDPR does not provide explicit standards for the structure or content of DDPs, leading to varying implementations of Article 15 across platforms [8].
Understandably, platforms that offer different types of services interpret and implement the GDPR provisions differently. For instance, the DDP of a digital marketplace like Amazon is naturally distinct from that of a digital public space like TikTok. However, platforms offering similar services should be able to provide comparable information to users. For example, platforms providing short-format video streaming services should, in principle, offer similar details through their DDPs. To the best of our knowledge, such a comparative study of the implementation of the right of access under the GDPR is rare in the literature. This observation leads to a key unexplored research question as follows: RQ1: How similarly do platforms offering similar services implement the right of access under the GDPR (Article 15(3))?
Desideratum 1: Comprehensibility: Irrespective of the answer to RQ1, these implementations (DDPs) should adhere to a set of requirements outlined in the GDPR to ensure user comprehensibility. According to Article 12(1), these DDPs should be (a) concise, (b) transparent, (c) intelligible, (d) easily accessible, and (e) using clear and plain language [12]. However, neither the GDPR nor accompanying guidelines [14] provide a clear technical definition of these criteria. Further, adherence to these requirements in the current implementations also remains unexplored in the research community. The authors in [8] came closest to understanding the question of adherence, but they restricted their analyses to conciseness. This gap leads us to the second research question in this paper: RQ2: How comprehensible are the DDPs provided by platforms under Articles 12 and 15 of the GDPR?
Desideratum 2: Reliability: Finally, another important expectation for the effective implementation of the provision is that the information in the DDPs is reliable. With the increased restriction on access to platform data for research [33, 4], many researchers are relying on data elicited through exercising the right of access for understanding the dynamic interplay between population and platforms. In these so-called “data donations” based studies, consenting participants first request their own data from the platforms, and then provide this data to researchers, typically for financial compensation. Several recent scientific studies use this setup to explore fundamental questions about digital platforms— such as user engagement [38], content recommendation [35], and ad targeting processes [36] — and even questions beyond online activities, e.g., offline risks [3] or the likelihood of addiction [37] using the resultant DDPs. However, if the reliability of the information provided by the platforms is compromised, then the insights drawn from such data sources might be invalid.
To this end, a recently conducted survey among researchers indicates that 86% of researchers report concerns regarding the reliability of the data collected using DDPs [34]. The most important dimensions of reliability that researchers are concerned about include: (a) completeness, (b) correctness, and (c) consistency of the shared information. This brings us to our final research question: RQ3: How reliable is the information within the DDPs provided by platforms under Article 15(3) of the GDPR?
The current work: Although the current work could have been conducted on any digital platform, due to the growing popularity in consumption of short–format videos [38, 25] and the impact they have on the population [13], we investigate the aforementioned research questions on Instagram, TikTok, and YouTube.  Figure 1 describes the pipeline of our conducted research. After analyzing the DDPs shared by these platforms, we conducted an extensive survey among 400 participants from Germany, France, Italy, and Spain to evaluate their assessment of the comprehensibility of the respective DDPs. Further, by utilizing (a) participants’ and (b) European Data Protection Board’s interpretations of the requirements, (c) current advancements in large language models (LLMs), and (d) our technical expertise, we propose data representations for the different categories of information. These recommended representations are then evaluated against the best representations of respective categories in the current implementation of the platforms by 200 participants.
To answer the questions regarding the correctness and completeness of the DDPs, we browse the three platforms using sock-puppet accounts to (a) automatically log the browsing behavior and (b) request the DDPs from each of the platforms. We repeat this process for a period of one month to understand the consistency in the shared DDPs for the accounts. In addition, to better understand the consistency across various accounts, we also collect DDPs requested by real-world participants from the three platforms.
Figure 1: Our pipeline to evaluate comprehensibility and reliability of the current evaluation of Article 15(3) of the GDPR.
We summarize our major findings with respect to each research question below.
RQ1: The studied platforms implement Article 15 of the GDPR differently: Instagram, TikTok, and Youtube share different categories of data in their DDPs. While Instagram and TikTok share more data with their users, YouTube’s shared data is rather limited. Irrespective of the amount of data, the only underlying similarity is that none of the platforms share the purpose of data collection and processing in their DDPs.
RQ2: Instagram’s current data representation is the most comprehensible: Based on the preference votes of 400 participants, we find Instagram’s DDP adheres to GDPR’s requirements more than the other two platforms.
RQ2: Considering interpretation from different stakeholders significantly improves the data representation: 200 survey participants prefer our proposed data representation to current representations along all requirements except conciseness. In fact, our proposal is preferred in as many as 44 out of the 52 evaluations (13 categories $\times$ 4 requirements).
RQ3: TikTok’s DDP is the most reliable as per our evaluation: Our analyses using sock-puppet accounts reveal TikTok’s DDP presents a complete, correct, and consistent representation of our browsing behavior. Instagram and YouTube exhibit varying degrees of missing data across DDPs collected over the period of our study. However, the analyses among real users show Instagram and TikTok share data for different periods in their DDPs for different categories, whereas YouTube shares all the data types for similar duration.
To the best of our knowledge, this study is the first detailed evaluation of adherence of different requirements mentioned in the GDPR. The insights drawn from our study further underline the need for standardization and follow-up on the implementations to ensure that the primary goals of data protection laws are achieved.
2 Background and Related Work
To position the current work in the perspective of the literature, we present a brief overview of the following: (a) GDPR and the rights of access, (b) leveraging data download packages for platform audits, (c) audits of implementations of provisions of the GDPR.
GDPR and the right of access: The GDPR is a landmark privacy law designed to give individuals greater control over their personal data and to standardize data protection rules across the European Union (EU) [12]. While the GDPR has many provisions and rights enlisted, our work focuses on Article 15(3): the right of access, which allows users to request a copy of their data collected by online platforms. Most platforms implement this requirement by providing a Data Download Package (DDP) upon request. For such implementations, GDPR provides a set of desiderata for the DDPs. Specifically, Article 12(1) requires a DDP to be concise, intelligible, transparent, accessible, and in clear and plain language [12]. Moreover, the GDPR gives the right to end-users to be able to port their data from one platform to another under Article 20. However, if platforms offering similar services implement these rights in significantly different ways, platforms will essentially deprive end users of their right to portability.
Leveraging data download packages: The richness of DDPs has enabled new research directions previously deemed infeasible. Indeed, many recent studies are utilizing them through donations made from participants with their explicit consent. To this extent, researchers have used DDPs to conduct research on topics related to personal health and safety [20, 37, 3], news and politics [6, 17, 18], auditing recommendation algorithms [24, 35], analyzing user behavior on social media platforms [38, 15], and understanding ad targeting types and user perceptions [36].
Audits of implementations of GDPR: Many studies have investigated (a) whether consent forms are designed properly by platforms [29, 30, 23, 9], (b) whether platforms are undermining user privacy under the pretense of ‘legitimate interest’ [31, 21, 22], (c) inconsistencies in platform’s data withdrawal behavior [11]. However, little attention has been given to platforms’ implementation of the right to access. A recent study shows that end users often view DDPs as overwhelming and unclear, leaving key questions unanswered while triggering privacy concerns [8]. Moreover, researchers have expressed concerns about the reliability of the DDPs and have faced challenges in using them [34, 19].
The current work: To the best of our knowledge, our work is one of the first studies that tries to unravel the current implementations of Article 15(3) of the GDPR from the lens of comprehensibility and reliability. Furthermore, our proposed recommendations for the potential implementation of DDPs may not only improve the current implementations, but also help in the implementation of other rights provided to users under the GDPR, such as the right of data portability.
3 RQ1: How similarly do platforms offering similar services implement GDPR Article 15?
Before analyzing the comprehensibility and reliability of DDPs, we first look into the content of the DDPs provided by each of the platforms (i.e., we compare how the platforms are currently implementing GDPR Article 15(3)).
3.1 Platforms under consideration
Since the primary focus of the study is to understand how similar platforms implement the same provision in the GDPR, we decided to focus on multiple platforms providing similar services. To this end, there are many options for digital streaming platforms (e.g., Netflix, Disney, etc.), digital marketplaces (e.g., Amazon, Zalando, etc.), and social networking platforms (e.g., Facebook, X, LinkedIn, etc.). In this study, we focus on three platforms that provide short-format video streaming services: TikTok, Instagram, and YouTube. Our motivation to study these platforms stems from the recent increase in popularity of short-format video platforms [38, 25]. Furthermore, all three platforms have been designated as Very Large Online Platforms (VLOPs) by the EU under the recently enacted Digital Services Act (DSA), indicating the far-reaching impact that these platforms have on the European society [13].
TikTok, Instagram, and YouTube implement Article 15(3) of the GDPR, through which they share DDPs with their end-users upon request. In the current implementations, platforms share data primarily in two formats (a) machine-readable – which is usually a single JSON or a collection of JSON files; (b) human-readable – where Instagram and YouTube opt for HTML version, whereas TikTok shares it in TXT files segregated across many directories. If one takes a closer look at the DDPs received from any of the platforms, their contents can be broadly divided into the following categories: (a) user’s usage, (b) user’s content, (c) personal details, (d) advertisements, (e) miscellaneous. Next, we examine the DDPs that Instagram, TikTok, and YouTube provide and observe their content and presentation. Most of the observations discussed below are summarized in Table 5 (see Appendix A).
3.2 Information regarding users’ usage
One of the most important facets of the information contained within the DDPs is how a user uses the social media platform. Such usage information may include the content that a user watches, searches for, or engages with implicitly or explicitly.
Watch history: Watch history is an ordered list of all the content that a user has watched before requesting the data. DDPs of all three platforms contain watch history data of the data subject. However, as Table 1 shows the data provided across the three platforms are widely varying. Among these, two points are strikingly different – (a) while YouTube provides this data for the entire lifetime of a user on the platform, Instagram provides data of at most the last two weeks and TikTok provides data for around six months before the data request (without any clear explanation for the reduced duration); (b) while TikTok and YouTube provide the watch history all at once, Instagram segregates it into three files based on the type of content, i.e., ads, posts, and videos watched by the user.
Feature YouTube Instagram TikTok
Duration Entire user’s lifetime 2 weeks 6 months
Content URL
Timestamp
Content title
Author name (with author’s URL)
Ad identification (whether ad or not)
File segmentation Single file
Three files
Single file
Table 1: Comparison of watch history data across platforms.
Like history: Similar to watch history, like history is an ordered list of all the contents that a user has liked before requesting the data. The details across the three platforms are noted in Table 2. The table demonstrates that there is a clear lack of consistency in how platforms currently implement Article 15(3) GDPR. While one platform (YouTube) records the like history on its application, it does not share this information in the YouTube DDP  [^1] Although Like history is not there in YouTube ddp, it is included in the Google DDP.. Another platform (Instagram) provides the watch history for just a couple of weeks without any link to the watched content, but it provides the like history of a lifetime, including references to the liked content.
Platform
Provides
like
history?
Details provided
TikTok
Link to the liked content,
timestamp of when it was liked
Instagram
Link to the liked content, author details,
timestamp of when it was liked
YouTube
Shows like history on mobile app and web
version but does not share it in the DDP
Table 2: Comparison of like history details across platforms.
Time spent on platform: While all three platforms have a detailed dashboard that mentions the average time spent and the exact amount of time users spend on the application daily, the same information is not provided in the DDP in response to Article 15 GDPR. On a positive note, TikTok recently added a new field in the data, namely “Activity Summary,” which enlists how many videos a user has commented on, shared, or watched until the end since their registration.
Other usage activities on platform:  Table 5 (in Appendix A) shows some other important usage activities. Most of the activities, e.g., comment, search, save, share, and writing a message are recorded by all platforms where the features are applicable. However, sharing across applications, i.e., when a user shares the video on other social media platforms or copies the link for posting elsewhere, is only shared in TikTok DDPs. Also, while both TikTok and Instagram maintain a list of inferred interests of the data subject, YouTube DDPs do not have them.
3.3 Information regarding user’s content
While the usage data is about how users consume or behave on video streaming platforms, information regarding their content refers to the content that a user creates and uploads to the platform for others to consume. This information encompasses the media (image/audio/video) a user uploads on their profile, along with any textual captions, locations, date and time information, etc.
Aspect Details provided Platforms
Content created
Media (image/audio/video),
textual captions,
date, and time information
TikTok,
Instagram,
YouTube
Addnl. details
Metadata such as software used
for uploading (e.g., Android gallery),
device ID, camera metadata
Instagram
Location details
Longitude and latitude of upload site
if the author tags the location
TikTok,
Instagram,
YouTube
Table 3: Details of user created content shared by platforms.
While all three platforms provide a copy of the media and text details, Instagram shares additional information including software used to upload the content, device ID, metadata about the camera, etc. (Table 3).
3.4 Information regarding personal details
While the two facets discussed above focus more on the content that users consume, engage with, or create, the information regarding personal details is more privacy-sensitive. These details contain basic personally identifiable information (PII), e.g., account details (which may include one’s name, phone number, e-mail address, date of birth, profile picture, etc.). However, apart from these basic personal details, most of the DDPs also include information about the different devices through which people visit the platforms. TikTok and Instagram directly store these details as part of the login information, whereas YouTube does not offer this data in its DDP. For YouTube, this data can be accessed only if someone requests their entire Google DDP. [^2] One justification may be that one can not sign in to their YouTube account unless they sign in to their Google account on the same device.
Login history: In light of the prior descriptions, login history contains the list of login activities made by the data subject. The login history details are significantly different in TikTok and Instagram DDPs. TikTok provides the timestamp, IP, device model, operating system, network type, and carrier provider in its DDP. On the other hand, Instagram’s login history data is very detailed. Along with providing the above details, Instagram provides (and therefore collects) cookie information, language code, Instagram app version, display properties of the device, hardware identifier, and some internal identifiers. Such differences in data collection and sharing practices point to two important possibilities: (a) TikTok might be collecting some of these details and not sharing them with its end-users and, therefore, is not being transparent; (b) Instagram might be collecting more information than necessary, thus becoming more intrusive to privacy. Diving deep into these possibilities is beyond the scope of this study, and we would like to explore these directions in future work.
3.5 Information regarding advertisements
Advertisement on social media platforms is a prevalent phenomenon, as users are constantly served with ads when they scroll through Instagram, TikTok, or YouTube. However, based on our observations, there is no standardized way of including ads that users have seen in their requested DDP.
Ads viewed: As mentioned in Section 3.2, Instagram maintains a separate list of ads viewed by users. However, the data provided by Instagram only mentions the author/advertiser name and timestamp. TikTok, on the other hand, does not demarcate any advertised content that a user has seen in their timeline. In contrast, YouTube distinguishes all the ads that a user has encountered from the organic content. Also, YouTube provides a link to the ad video along with its title and when it was viewed. However, neither of the three platforms provides information on ad targeting parameters. From the ads viewed details, a user can understand these are the ads they got to watch on their timeline, but why those particular ads were served to them remains an enigma. While this information does appear when an advertisement or paid content appears on the three platforms, it is not included in the shared DDPs.
Ad targeting data sources: The Instagram DDP contains a list of advertisers who have used data about one’s online activity or profile for targeted advertising purposes. This list also conveys information about how these advertisers obtained users’ data. In specific, as per this information, there are three ways in which the advertiser might target a person: (a) custom audience – where the advertiser targets the data subject through a custom audience with a list of customer data (e.g., email addresses, phone numbers, or other identifiers) which the advertiser might have obtained from other third party sources; (b) remarketing custom audience – where the advertiser targets the data subject because (s)he might have interacted with some of the advertiser’s content, visited their website, etc.; (c)  in-person store visit – where advertiser can use location data or check in data because the data subject might have visited one of their physical stores. The DDPs of YouTube and TikTok do not provide any such information on advertisers and how they gather data about a particular user.
Off-platform activities: Another important source of data through which these platforms collect enormous amounts of digital behavioral data is off-platform activities. This is a list of all the activities in external platforms (outside the source platform) that these platforms have tracked or linked to a user. Note that this might be an artifact of the said external platforms’ installed Meta/Google/TikTok pixels. A user might have used a Meta or Google login to sign up, other tools embedded in the software development kit of the external platform, or any data-sharing agreement between the external company and any of the platforms we investigate.
For instance, if a user visits the website or application that also installs one of the tracking techniques of any of the short-format platforms, then the said website can share the user’s activities, such as page view, purchase, search, logins, etc., along with other activities, such as adding a product to the cart, view contents, install app, launch app, etc. Instagram and TikTok DDPs contain off-platform activities of users, whereas YouTube does not have such data. Both Instagram and TikTok share the external platform name, event or activity type, and the timestamp in their DDPs.
The DDPs from the three platforms also provide a wide variety of other details. For example, they may contain information on link history, survey, application setting, shopping, etc. For more details on this, please refer to the summary of all the discussions above, listed in Table 5 (see Appendix A). All three platforms miss one of the important goals set by GDPR Article 15(1)(a), which is to include the purposes of the data processing and collection. None of the DDPs include a purpose statement explaining why certain data is collected, nor do they provide a README or manual to help users understand the significance of the data being collected.
Important takeaways
[…]
  The three platforms implement Article 15 very differently. The Instagram and TikTok DDPs contain a significantly larger number of data types than the YouTube DDP. Instagram shares (and thus collects) more data.
[…]
4 RQ2(a): How comprehensible are DDPs?
Here we investigate whether the current implementations of platforms adhere to the requirements Chapter 3 of the GDPR outlines a number of rights of data subjects (i.e., end users), including the right of access (Article 15) [12]. Before outlining these rights, the GDPR provides a list of requirements for how information should be communicated to end users. Specifically, the first paragraph of Article 12 of GDPR reads, “The controller shall take appropriate measures to provide any information referred to in Articles 13 and 14 and any communication under Article 15 to 22 and 34 relating to processing to the data subject in a concise, transparent, intelligible and easily accessible form, using clear and plain language $\dots$ [12].
As the GDPR does not provide more succinct definitions of these requirements, we adopt two approaches to interpret these terms. First, following trends in empirical contract research to employ user surveys for interpretation of legal terms [5], we conduct a large-scale user survey among participants from different European countries to understand their interpretations of some of the requirements. Second, we apply the interpretations of the requirements by the European Data Protection Board (EDPB) [14]. Finally, we evaluate the current implementations, i.e., the DDPs obtained from Instagram, TikTok, and YouTube, against the interpretations by users and the EDPB.
Participant recruitment: We recruited participants to conduct our survey from the crowd-sourcing platform Prolific[27]. On Prolific, we hired 100 participants each from Germany, France, Spain, and Italy (400 in total) who had a high approval rate of at least 98% to take part in our survey. Our motivation for a general set of respondents stems from the fact that the GDPR entitles general residents in the European Union(EU) with the right to data access, and requirements are intended to benefit these residents. Though, eventually, our plan is to extend the study to all countries within the EU, we chose to start with the countries that have the highest GDP [32]. We also selected the standard gender breakdown while setting up the surveys. In total, we recruited 228 males (57.0%), 161 females (40.3%), 9 self-reported as “other” (2.3%), and 2 participants who preferred not to disclose their gender (0.5%) (See Table 6 in Appendix C for details).
Prior to entering our survey, we presented an online consent form for expressing explicit consent from the participants. Upon completion of the survey, we compensated each participant at a rate of £9 per hour, which is recommended by Prolific to be a good and ethical rate of remuneration [26]. Our survey comprised three components: (a) Awareness of participants about the rights, (b) Interpretation of the requirements by the participants and (c) adherence evaluation. Next, we elaborate abut the survey setup and important observations for each of the components.
4.1 Awareness about the GDPR right of access
Survey setup: In the first component, we first tried to educate our participants about Article 15 of the GDPR and how they can request their data from online platforms. We also provided a link to a Google Drive folder where they could see an example of a DDP. At this stage, we asked our participants three questions regarding (1) whether they were aware of such right before taking our survey, (2) whether they had exercised this right by requesting their data, and (3) if they had exercised their rights, then why. We designed this part of the survey to understand the general awareness of the participants and make them aware of the said right.
Observations: Out of the 400 participants, nearly $72\%$ participants responded that they were aware of the right of access to data before participating in our survey. At the same time, only $29.2\%$ of them answered in affirmation when it comes to exercising their rights by requesting the data on some platform. Although there is a massive gap between awareness and exercise of the data requests, the numbers are surprisingly high. We also asked our participants to provide reasons for exercising their data requests in free-form tex. Through manual annotations the responses were characterized into four groups. Almost half of the participants (51%) mentioned curiosity, knowing what information platforms collect about them, to be the primary reason for their data request. Further, 18% mentioned seeking specific information, e.g., identifying a song they had listened to earlier, cross-checking specific details, or determining the amount of time spent on the platform, etc., as the primary reason for their requests. Finally, 11% of our participants made their requests to keep a backup.
Interestingly, one-fifth of the participants (20%) mentioned that they requested data from several platforms for participating in some research study. Note that several recent studies–for understanding social media– use data donations from users on Prolific ([37], [35]) and other crowd-sourcing platforms as their primary data source. This observation likely explains the surprisingly high percentage of awareness and exercise statistics that we reported above.
4.2 Interpretation of the requirements
Survey setup: To understand the interpretation of requirements by common end users, in the second part of our survey, we showed the participants the paragraph of Article 12 (see above), which mentions the different requirements of DDPs. We asked our participants about their interpretation or expectation of the shared personal data in the DDPs to follow any of these requirements. Next, we elaborate on the different interpretations of these requirements from the responses of our participants while contextualizing them further with the guidelines of the EDPB [14], which represents the data protection authorities of all EU member states.
Conciseness Interpretations by the survey participants: We perform thematic analysis of the responses provided by the 400 participants using BERTopic [16]. The important themes that resonated across the responses spanning all countries are – prioritize important & relevant data, summarize important observations, and present data in a structured way that is easier to navigate.
However, a fourth theme emerged in the interpretations shared by participants from Italy, France, and Spain, but not by those from Germany. Specifically, these participants touched upon the concept of data minimization. To this end, some participants mentioned the documents should be shorter, while others mentioned avoiding an overwhelming amount of text. In short, while participants from Germany only prioritize the informativeness of the content, other participants prioritize the length of the content as well.
Interpretations by the EDPB: In this context, the European Data Protection Board interprets concise presentation to be done efficiently and succinctly to avoid information fatigue [14]. While the above interpretation falls within the interpretations mentioned by our survey participants, the EDPB also asserts that the presentation should differentiate privacy-sensitive information from non-privacy-sensitive information.
Transparency Interpretation by the survey participants: Unlike the case of conciseness, there are no major differences in how participants from all countries interpret transparency. The interpretation of transparency can be divided into two major aspects. First, participants call for unambiguous, full disclosure of all the data that has been collected about them in the way they are collected. Second, and most importantly, some participants also expect the DDPs to mention how the data is processed and for what purpose the data is collected. To this extent, the response from one of the participants captures all the notions beautifully: “It (Transparency) means providing clear honest and complete information about how the data is being used. There should be no hidden details or surprises for the individual. People should easily understand what data is being collected why its being collected who will access it and how long it will be kept.” At this point, we would also like to note that as per the participants’ interpretation, conciseness and transparency are two competing requirements, i.e., to satisfy one, the other possibly has to be violated.
Interpretation by the EDPB: While the EDPB draws most of its interpretations from Recital 39 of the GDPR, it interprets transparency as allowing data subjects to determine the scope and consequences of the concerned data processing [14]. To this end, the EDPB calls for data controllers (i.e., platforms) to separately spell out the most important consequences and risks that the data processing may entail.
Intelligibility Interpretation by the survey participants: Among all the requirements provided by the GDPR, the most subjective one is that of intelligibility. However, almost all the participants explicitly interpret it to be easy to understand and link it closely with the language being straightforward and without any technical and/or legal jargon. Participants from all countries also mentioned that intelligible data should have a clear explanation of how and why a certain piece of information is collected. The participants also noted some interesting expectations in this regard. For instance, participants from Germany mentioned that platforms shall share their DDPs in one’s language of comfort (native language or other language of choice). Moreover, participants from Italy stated that platforms should share their DDPs in the form of elegant visualization.
Interpretation by the EDPB: The EDPB describes intelligibility to be a representation that should be understood by an average member of the intended audience. In addition, they also mention that it is closely linked to the requirement to use clear and plain language.
Clear and plain language Interpretation by the survey participants: Understandably, most of the participants interpret plain and simple language to be words that are easily understandable and unambiguous. Some participants elaborate on the nuance by stating that legal and technical terms should be avoided or at least explained in a simpler way. To this end, one clear observation from our survey responses is that participants do not distinguish between intelligibility and language.
Interpretation by the EDPB: Along the lines of the interpretations mentioned above, the EDPB interprets this requirement as being definitive and should not be phrased in ambivalent terms or leave room for different interpretations [14].
Accessibility Interpretation by the survey participants: Most of our participants interpret accessibility as being easily available, i.e., it is easy to find the place to request the data and to access the data. However, there are certain interesting and valid interpretations which are worth mentioning. Some participants mentioned that the data needs to be available in a format so that it can be opened without any sophisticated software or tools. At the same time, the timeliness at which the data is provided upon request is another interesting aspect touched upon by some of our participants. Another interesting interpretation was the data should be available to everyone, including people with visual impairments by using accessible design features like screen readers or alternative text making sure the users can access data without any barriers.
Interpretation by the EDPB: The EDPB’s interpretation of accessibility also follows from the above popular responses from the survey participants, wherein they say that it should be immediately apparent to end-users as to where and how the DDP can be accessed [14].
Our observations on accessibility: Despite agreeing with all of the above interpretations, we observe a few differences across the platforms because accessibility is the most objective property for evaluation. In terms of ease of request, from the content page of a user’s account, one needs to click 6, 6, and 10 times on TikTok, YouTube, and Instagram, respectively, to be able to request the data in the human-readable format. Notice that, while on Instagram, a user needs to click more, the platform provides many options to the end-user, such as the time duration that their DDP should contain. Moreover, based on our anecdotal observations, the turnaround time for TikTok upon request was instantaneous during this research study. For Instagram and YouTube, it often takes 10 to 15 minutes for the DDP to become available. However, all the turnaround times are within the GDPR-prescribed maximum duration of one month. Further, one can download one’s data on TikTok and Instagram from both their applications and web versions, but YouTube shares the DDP with its end users by sending an e-mail with a URL to the DDP. All these platforms provide DDPs in a compressed format, which makes it harder for the end-users to open them on a smartphone. At the same time, an end user needs to have a browser or text reader to open the HTML (YouTube and Instagram) and TXT (TikTok) DDPs. In summary, while the data request and download are relatively accessible in the current setting, reading the data is still complicated for the average end user. At this point, we would also like to note that we are not aware of any special provisions that platforms may have for people with visual impairments to access the data.
4.3 Adherence evaluation
Survey setup: In the final part of the survey, we presented our participants with the information shared in the DDPs of Instagram, TikTok, and YouTube in an anonymized and randomized format. We showed different categories of content and the whole DDP to our participants. The different categories of content that we shared include data from all of the categories of information that we elaborated in Section 3.  Table 7 (refer Section C.1) summarizes the specific thirteen categories that we evaluated. [^3] We should point out that YouTube does not provide like history, autofill, location, and off-platform activity. Hence, for these categories, we asked our participants to evaluate between Instagram and TikTok. Note that, for this comparative study, we used DDPs that are shared in a human-readable format by all three platforms. Instagram and Tiktok share the human-readable version in HTML and TXT formats, respectively. However, YouTube DDP only contains watch history, search history, personal information, and login history in HTML format. All other categories of data are found in CSV or JSON formats. Since the primary goal of this study is to compare the contents, we showed them consistently in a $<category,value>$ pair in our survey. We acknowledge that the way data is shown to users may affect some of the results. However, in this study, we restrict ourselves to the evaluation of what content is shared rather than how it is shown. We plan to explore the visualization aspects in future studies.
For each category, we showed participants the content shared by the three platforms side by side and asked them which platform’s data representation they found to be more concise/intelligible/transparent/in clear and plain language. [^4] We exclude accessibility from the evaluation because it pertains to getting access to request and download DDPs more than to their representations. Also, we stated our observations about this requirement in Section 4.2.  Figure 8 (see Section C.1) depicts a representative example of one of the questions that we used in our survey. The first preference in Figure 8 means the most preferred choice of data representation of a participant, whereas the third preference means their least preferred choice of data representation. We posed this question as a multiple-choice grid, so each participant had to provide an ordered preference for the three representations. For example, if a participant rates Instagram’s representation as their first preference, then they cannot rate any of the other platforms as their first preference.
Note that the final part of the survey consisted of 14 (13 categories + overall DDP) $\times$4 evaluations where participants gave us an ordered preference. Hence, to avoid participant fatigue, we asked 50% of our participants to evaluate intelligibility and transparency and asked the other half to evaluate conciseness and language properties. Hence, 200 participants answered each of the evaluation questions.
Observations
Next, we report our findings on the adherence evaluation of four requirements.
Language: Figure 2 shows the percentage of first preference votes elicited by the representation of each of the platforms for the four requirements that we surveyed. In the evaluation of clarity and simplicity of language, Instagram won the first preference votes in eleven out of the thirteen data categories (Figure 2(a)). In contrast, for the remaining two categories, participants preferred the language in which YouTube presents its search and watch histories.
To understand the overall preference of participants across the board, we weight every first preference vote with a score of 3, second preference vote with a score of 2, and third preference vote with a score of 1. For categories, where we are comparing only two platforms, we assigned scores 3, 1 for first, second preference respectively. Then, we evaluate an average preference score for each of the platforms based on participants’ preference orderings. Figure 3 shows the average preference scores of each of the platforms in each category of data. Based on the average preference scores, apart from watch, search histories and location, the gap between preference toward the language of Instagram’s data representation is way more than that of TikTok and YouTube (Figure 3(a)). Across the different categories of data, the participants’ average preference score for Instagram’s language is 2.48, whereas TikTok and YouTube fail to reach an average preference score of 2.
Apart from the thirteen categories of data, we also asked our participants to evaluate the overall DDP from the three platforms. When participants were asked to evaluate the overall DDP, 55% of them preferred the language of Instagram’s DDP over the other two. TikTok and YouTube DDPs are preferred by 25% and 20% of the participants, respectively.
Intelligibility: Similar to Language, Instagram won the first preference vote for eleven out of the thirteen categories evaluated (Figure 2(b)). Among these, except for devices and location categories, nearly more than 75% of the participants preferred Instagram’s data representation to be the most intelligible.
The average preference scores almost follow the same trend as that of language, where the Instagram DDP’s intelligibility (2.53) is substantially higher than that of YouTube (1.75) and TikTok’s (1.64) (Figure 3(b)). For the evaluation of the overall DDP, 63% of the participants identified the Instagram DDP to be the most intelligible. TikTok and YouTube were chosen as the first preference by a considerably smaller proportion of our participants, nearly 18% each.
Transparency: YouTube won the first preference for five categories, where over 70% of the participants chose it for watch history, search history, and login, followed by devices and content with 55% and 46%, respectively (Figure 2(c)). For the remaining eight out of the thirteen categories, Instagram won the first preference of the participants in a large percentages.
We find that the average preference scores of Instagram and YouTube are comparable (Instagram: 2.23, YouTube: 2.15), whereas TikTok (1.66) has lower score (Figure 3(c)). In the evaluation of the overall DDP, Instagram is preferred more compared to the other two platforms, with nearly 46% of the participants choosing it to be the most transparent. On the other hand, YouTube is preferred by 35% of the participants, and the remaining 19% selected TikTok.
Conciseness: Based on Figure 2(d), we found that in nine out of thirteen categories of data that we surveyed, respondents preferred TikTok’s data representation as the most concise, whereas Instagram’s data representation is found to be the most preferred for the remaining four categories. While TikTok’s representation is preferred substantially in the location, autofill, and off-platform categories ($\geq$60first preference votes), Instagram’s representation is preferred more for likes history and contents ($\geq$65first preference votes). Based on the average preference scores for conciseness, we see that the representations are found to be very comparable in most of the data categories (Figure 3(d)). The average preference scores of Instagram and TikTok are very comparable (Instagram: 2.12, TikTok: 2.15), whereas that of YouTube is considerably lower (1.60). While evaluating the overall DDP, 43% of the participants found Instagram’s DDP to be more concise, followed by Tiktok and YouTube with 35% and 22%, respectively.
Variations across countries: We do not observe qualitative variation across the four countries for three requirements – language, intelligibility, and transparency. However, for conciseness, participants from Germany predominantly preferred Instagram, with over 50% of the participants selecting it as their first choice across most categories. In contrast, participants from other countries primarily selected TikTok as having the most concise representation for all the categories, except likes, connections, and content. To understand this, we manually examined how participants across different demographics interpret conciseness. We found that, except for participants from Germany, most demographics have an interpretation of conciseness as: (1) minimizing document length, (2) avoiding information overload etc. This difference in interpretation is reflected in how our participants evaluated the data representations: while participants from Germany gave more priority to content, other participants gave more priority to data categories having less content.
(a) Clear and Plain Language
(b) Intelligible
(c) Transparency
(d) Conciseness
Figure 2: Percentage of first preferences across all categories and for the entire DDP. Instagram is the top choice for most categories across all requirements, except for conciseness. For watch and search histories, YouTube is the most preferred platform in transparency, intelligibility, and language.
(a) Clear and Plain Language
(b) Intelligible
(c) Transparency
(d) Conciseness
Figure 3: Comparison of 13 different data categories, including an evaluation of overall DDPs for the four requirements. Instagram achieves the highest overall DDP score for all the four requirements.
Important takeaways
[…]
  We found participants’ interpretation of the requirements are in line with that of European Data Protection Board. However, there seems to be conflicting interpretations of conciseness and transparency among both stakeholders.
[…]
5 RQ2(b): Improving DDPs comprehensibility
Figure 4: Comparison of watch history representation in our proposal and that of YouTube’s current DDP implementation.
(a) Clear and Plain Language
(b) Intelligible
(c) Transparency
(d) Conciseness
Figure 5: Percentage of votes for the proposal and the earlier winning platform across different categories for each requirement. Our proposal emerged as the top choice in all fields for the language and intelligible requirements.
In Section 4, we observed that none of the platforms adhere to the prescribed requirements of the GDPR properly across all data categories, highlighting the need for better representations. Motivated by this, here we utilize the different interpretations we elicit from survey participants to come up with novel recommendations that lead to better representations.
5.1 Methodology
To come up with an actionable recommendation for data representation, we take the following steps.
Step 1: Summarize the interpretations: As a first step, for each of the requirements, we collect the interpretations from our surveyed participants and that of the EDPB [14]. Next, we feed this entire set of 401 interpretations to a large language model (LLM) – Gemini 2.0 Flash [10]– using the Google AI Studio platform. We ask the LLM to generate a set of recommendations that encapsulate the most prevalent interpretations by summarizing the above responses. For reproducibility purposes, we make the prompts we used and the responses in  Appendix B. Note that in this approach, we include the interpretations of not only the data subjects, but also of data protection authorities (through the interpretation of EDPB).
Step 2: Generate recommended data field for each data category: In the second step, by providing the top recommendations for each of the requirement, we ask the LLM to make some assumptions. Specifically, we state that we are data engineers working in a short-format video platform operating in the European Union. Moreover, an end user requests us to provide them with their personal data for a given data category. Then, we ask the LLM to let us know of the specific data fields that we should provide to the end user. Subsequently, the LLM provides us with a set of recommendations for data fields of information in the queried category in the form of a JSON object (available in Appendix B).
Step 3: Manual inspection by researchers: In the final stage, three researchers who are part of this study and are conversant with DDPs of a whole array of platforms assess these recommendations. The primary goal of their assessment is to remove any instance of hallucination in the form of potentially irrelevant data fields that may have been recommended by the LLM. Next, given their expertise in understanding data donations in other platforms like Netflix, Prime video etc., we ask them to suggest any additions or removals to the recommended fields. By taking these steps, we ultimately come up with the final proposal of recommendations. Note that while Step 1 introduces the end users’ and data protection authorities’ interpretations, this step also instils the common interpretations of tech experts on other platforms and researchers who want to unravel the dynamics of platforms using these DDPs.
An example of the recommended representation for watch history and that of the most adhering representation among the existing one’s (YouTube’s) is shown in Figure 4. Note that apart from the details about the video watched, the proposal also shares the duration of the video and what fraction of it was watched by the user and on which device and network.
5.2 Evaluation of the proposed representation
To evaluate our proposed representation, we conducted another survey among 200 participants (who have not participated in the earlier survey mentioned in Section 4) from Germany, France, Spain, and Italy. We follow the identical participant recruitment strategy as elaborated in Section 4. Table 8 (See Appendix D) shows the user demographics. During the survey, we showed the participants two representations for the evaluated categories: (a) proposed representation: which we generate from the steps mentioned in Section 5.1, (b) winner representation: for each data category, the representation among Instagram/TikTok/YouTube which won the first preference along more requirements (i.e., the representation which participants suggested to be more adhering to the GDPR). We anonymize and randomly present each pair of representations to the survey participants. Then, we ask the participants to select one of the two representations that they find to be more intelligible/transparent/concise/in clear and plain language.
Observations We compare the representations of the data category fields of our proposal and that of the best among the existing short-format video platforms for 13 different categories across four GDPR requirements. [^5] Notice that the data categories are slightly different than the evaluations done in Section 4.3. We are not considering personal information, as there were no suggestions for adding or removing any fields. Instead, we include activity summary, which is provided only by TikTok, for comparison.  Figure 5 shows the percentage of votes for the two representations obtained in the 52 evaluations (13 data categories $\times$ 4 requirements), that we surveyed. We observe that the participants preferred our proposed representation in as many as 44 categories. However, their preferences vary for different requirements.
Language and intelligibility: For clarity and simplicity of language and intelligibility of the representation, our proposal gets upwards of 67% of the votes in most of the categories (except language in comments and intelligibility in autofill categories) suggesting our proposed representation is both intelligible and in clear and plain language.
Transparency: Our proposal is voted to be more transparent than the best of the existing short-format video platforms in 11 out of the 13 categories. Apart from search history, in all the other winning 10 categories, our representation gets upwards of 75% of the votes. In watch history, the split between our proposal and that of YouTube’s DDP is 47% and 53%, respectively, indicating that both representations are of comparable quality. However, in autofill information category, Instagram’s data representation gets nearly 93% of the votes from our survey participants. However, we argue that having less autofill information is indicative of less privacy-intrusive data collection. While our proposed recommendation includes only the email address, name, address, and phone number, that of Instagram contains much more sensitive information.
Conciseness: Notice that in Section 4.2, we mentioned conciseness and transparency, and to some extent intelligibility, are mutually competing requirements. We observe the same conflict in the results of this survey. Unsurprisingly, our proposal was preferred to be the most concise representation in merely 7 out of the 13 categories. Even then, in watch history, search history, login history, and device details – our proposed representation was found to be more concise by upwards of 75% of the participants. It is also worth mentioning that, much like the survey presented in Section 4.3, the participants from Germany have different preferences compared to those from other countries. Except for autofill information, over 66% of German participants preferred our proposed representation in all other categories.
Despite our proposed recommendation coming out as a significant improvement over any of the existing data representations, achieving this was not the main objective of this section. Our primary intention of this section has been to convey how, relatively easily, one can come up with more effective data representations by taking into account the interpretations of different stakeholders. This potentially indicates that currently, platforms are probably operating in silos without much deliberation and sincere effort into implementing Article 15 of the GDPR. At the same time, because of the mutually conflicting nature of the requirements, it is harder than anticipated to satisfy all the requirements with a single representation.
Important takeaways
[…]
  Taking into account the interpretations of different stakeholders (e.g., end users, legal scholars, technical experts, and researchers) can offer a significant improvement to the data representations across different categories.
[…]
6 RQ3: How reliable are the DDPs?
In this section, we investigate how reliable the DDPs are. We first define the requirements for quantifying the reliability of DDPs. Next, we outline the data collection employed to evaluate their reliability. Finally, we conduct a comprehensive assessment of the reliability of DDPs for the three platforms.
6.1 Requirements for reliability
86% of researchers who participated in a Data Donation Symposium reported encountering problems with the DDPs provided by various platforms [34]. The primary concerns identified include: (1) incompleteness, such as missing data types or limited time ranges, and (2) inaccuracy, such as wrong timestamps, incorrect ordering which renders the data unreliable to both end users and researchers. Therefore, due to the popularity of using DDPs to conduct research, there is a pressing need to systematically assess how reliable DDPs are. Hence we evaluate the reliability of DDPs on three requirements.
Completeness: Evaluates whether the platforms provide a comprehensive record of all user activities or whether the data is subject to sampling or omissions.
Correctness: Determines whether the sequence of recorded activities in the DDP accurately reflects the chronological order of events and whether any wrong entries are present.
Consistency: Assesses the uniformity of data across users and within users across different retrieval instances.
6.2 Experimental setup
We collect two types of data: (a) from real-world users, and (b) utilizing bot accounts.
Real-world users Inspired by prior work, we designed our own data donation website, which is privacy preserving and allowed us to obtain DDPs from real-world users [38]. To ensure user privacy, all collected data was anonymized and any PII was discarded on the front-end of our data donation website. During the data collection process, we intentionally mandated the donation of specific data categories for each platform. For TikTok and YouTube, video browsing history was mandatory. In the case of Instagram, the mandatory categories included (a) like history and (b) ads and topics, which consist of posts, ads, videos viewed, and ads clicked. Other data categories, such as search history, comments, etc., were optional and users provided them at their discretion. Private messages and other sensitive data were not collected. Users were compensated based on the categories and amount of data they donated. The maximum payment offered was \$16 for TikTok, \$9 for YouTube, and \$14 for Instagram data donations. Data collection for TikTok and YouTube took place between August and September 2024, while for Instagram, it was conducted between December 2024 and January 2025. We collected data from the same set of 33 users across all three platforms. 17 (51.5%) male and 16 (48.5%) female participants donated their data.
Bot accounts We automated the process of video browsing and video liking activities using pyautogui [2] to control mouse and keyboard and opencv-python [1] to detect the like button on the desktop screen. This approach differs from traditional browser automation [7], as we directly controlled the desktop environment instead of interacting with the browser via scripts. We captured interactions with the browser by downloading the HAR (HTTP Archive) files (HAR is a JSON-formatted file that logs all network traffic between the browser and the server). We viewed between 20 to 25 short-format videos in each video browsing session on each of the three platforms. Our bot accounts watched each video for a random duration, ranging from 15 to 60 seconds, and they randomly liked some of the videos. This approach allowed us to gather accurate ground truth data regarding user activities. The process was conducted for nearly thirty video browsing sessions between October 2024 and December 2024 for each platform. To conduct our analysis, we collected multiple DDP snapshots from each of these bot accounts.
6.3 Observations
Completeness: As mentioned in Section 3.2, Instagram and TikTok do not provide the entire watch history of a user, since they register on the platform, without any explanations. Hence, these DDPs are inherently incomplete and potentially violate Article 15(1)(d). However, with whatever data they share, to evaluate completeness, we utilize the data collected from the bot accounts. To this end, we compared the number of activities recorded in the HAR dump with those in the DDP for each video browsing session. We used the DDP that was closest in time to the corresponding HAR dump. Instagram offers two video browsing feeds: (1) home feed, and (2) reels feed. Interestingly, we found that videos watched under the reels feed were not present in the DDP, although like activities associated with them were recorded. For Instagram’s home feed, the percentage of recorded entries in the DDP for browsing videos and likes is 100%. On TikTok, all browsing and like activities were fully recorded, with a completeness rate of 100%. For YouTube, the recorded entries accounted for 99.5% of the activities observed in the HAR dump.
Correctness: After assessing the completeness of DDPs, we evaluated its correctness using the same data from bot accounts, by checking whether the order of contexts was maintained and whether any arbitrary entries were present in the DDP or not. To achieve this, we measured the Jaccard similarity for the following aspects between HAR and DDP for each activity type.
$\bullet$ Date: Consistency of recorded timestamps. We compared them up to the minute level to allow for slight variations.
$\bullet$ Context: Consistency of video IDs or author IDs.
$\bullet$ Overall: A combined evaluation of date and context. Table 4 presents the average Jaccard scores across the video browsing sessions for the three platforms. For Instagram, there is a drop in the Jaccard scores of browsing, because of changes in usernames, i.e, account username change or account deletion. For TikTok, Jaccard scores achieved 100% accuracy across both browsing and like activities. For YouTube, the average Jaccard scores are 100% for date, 99.83% for context, and 99.5% overall. The slight drop in the overall score is because of the 0.5% missing entries discussed earlier. For TikTok and YouTube, the recorded timestamps showed a minimal difference of $\pm$ 5 seconds compared to the HAR dump values for the same activity. However, for Instagram, the difference was as high as up to one minute, leading to a reduction in the accuracy for date and overall.
Platform (Activity) Date Context Overall
Instagram (Likes) 100% 100% 100%
Instagram (Browse) 96% 97% 91%
TikTok (Likes and Browse) 100% 100% 100%
YouTube (Browse) 100% 99.83% 99.5%
Table 4: Average Jaccard scores across different video browsing sessions for the three platforms. Concerning correctness, TikTok sits at the top, followed by YouTube and Instagram.
(a) Instagram
(b) TikTok
(c) YouTube
Figure 6: Ratio of entries retained across multiple snapshots for the same account: Instagram (like history), TikTok (video browsing history and like activities), and YouTube (video browsing history).
Consistency (across snapshots): Next, we determine whether a particular account receives consistent data when making multiple requests from the same platform. Towards this goal, we analyzed differences between multiple DDPs obtained from bot accounts. We examined whether the entries present in an earlier snapshot were retained in a later snapshot. For Instagram and TikTok, the snapshots are collected with a gap of 1 week. For YouTube, we use a gap of one month.
Figure 6 presents the ratio of overlapping entries between two DDPs to the total number of entries in the earlier DDP for three key aspects, i.e., date, context, and overall. For Instagram, despite using the same bot accounts, we observed variations in the covered duration of the video browsing history across multiple snapshots, ranging from 6 to 13 days. Hence, due to limited available video browsing history, we solely focus on analyzing the like history.  Figure 6(a) shows the results for Instagram. We identify some missing entries in the later snapshots. Specifically, 6% of the entries are found to be missing from snapshot 3 to snapshot 4.  Figure 6(b) illustrate the results for video browsing history and like activities on TikTok. Based on our findings, we observe that TikTok provides consistent data across multiple retrieval attempts. Lastly, in  Figure 6(c), we observe that approximately 12% of entries were missing for snapshot pairs (1, 3) for YouTube. Among these, nearly 62% entries were advertisements, while the other videos got deleted from the platform itself.
Consistency (across users): To assess whether all users receive a similar amount of data, we analyzed video browsing history, search history, and like history obtained from real-world users across the three platforms. First, we measured the duration of data that was contained in each activity history by calculating the time difference between the earliest and the latest entries in the respective activity history lists. For Instagram, we defined browsing history as the combination of ads viewed, posts viewed, and videos watched, while we included both keyword or phrase searches and user searches in the search history. To understand the extent to which the distribution of provided data to the users differs, we analyzed the cumulative distribution function (CDF) of the durations of video browsing, search, and like history across the platforms.
Figure 7 presents a comparative view of activity durations across users. For Instagram, we observe distinct clusters of users with browsing history durations concentrated around 6 and 13 days (Figure 7(a)). This non-uniform distribution indicates that some Instagram end users receive significantly different amounts of data. Similarly, on TikTok (Figure 7(c)), video browsing and search history durations exhibit substantial variation. Specifically, some end-users had data spanning approximately 180 days, while others had nearly 450 days of recorded video browsing and search activities. To ensure a fair comparison, we specifically restricted our analysis to TikTok users whose earliest recorded activity occurred before March 2023. This criterion allows us to focus on the subset of users who have had ample time to accumulate browsing and search history. As a result, we analyzed data from 23 TikTok users. These findings highlight TikTok’s tendency to provide data inconsistently among its users. However, such kind of clusters were not observed for like history on both Instagram and TikTok. The analysis of video browsing and search history on YouTube (Figure 7(d)) does not reveal any discernible patterns, suggesting that the platform may provide DDPs more consistently to its end users.
(a) Instagram - Browse, Search
(b) Instagram - Likes
(c) TikTok
(d) YouTube
Figure 7: Plots illustrating the CDF of data duration provided to users for three activities – search, browse, and like history. On Instagram, clusters in browsing history are observed at 6 and 13 days, while on TikTok, clusters appear at 180 and 455 days for both search and browse history.
Important takeaways
[…]
  In terms of correctness and completeness, TikTok provides the most reliable DDPs. Also, its DDPs are consistent across multiple retrievals.
[…]
7 Concluding discussion
Summary of the insights: To the best of our knowledge, ours is the first systematic attempt to evaluate the implementation of the right of access (Article 15) across platforms offering similar services. Our detailed evaluation of Instagram, TikTok, and YouTube, focusing on their DDPs, reveals that Instagram and TikTok share more data categories, while YouTube’s shared data is limited. The only underlying similarity that cuts across platforms is that of non-disclosing the purpose of data collection and processing. While Instagram’s data representation is deemed most comprehensible, TikTok’s is found to be the most reliable. Finally, a proposed data representation, leveraging interpretations of different important stakeholders, is found to be preferred by participants across multiple countries along the different requirements.
Limitations of the current work: Like any other work it has its own share of limitations. The survey population could have been more diverse and representative of common users on the street. However, note that our work is comparative by nature and we do not expect that the choice of participants will impact the relative rankings and comparisons presented across the paper. Further, our study solely focuses on data shared by platforms upon request and not the data that is actually collected from the users. Including this information may further enrich the quality of the work.
Recommendations for stakeholders: Platforms should improve the consistency among different categories of data that they share with end-users. The way different related categories of data are being shared differently (e.g., like vs. browsing history on Instagram) raises concerns about the lack of coordination within the platform while implementing the regulations. The other significant room for improvement for platforms is increasing the comprehensibility of the shared data. As we show in our work if we consider the interpretation of different stakeholders, the resulting representations can already be a significant improvement on the current representations.
The current differential and ineffective implementation not only affects end-users’ rights of access but also affects their rights to port their data (Article 20). Such rights of users will only be useful if data protection authorities provide more detailed technical specifications and/or standardization for implementation. As Section 5 shows involving different stakeholders may further improve the standard of implementations. Moreover, authorities need to enforce these requirements–otherwise platforms may not have any ramifications to worry about if they do not comply with them. To this end, the current study also indicates how effective enforcement of the GDPR rights may benefit from computer science research by relying on audit strategies to observe compliance at scale.
Acknowledgments
We would like to extend our sincere gratitude to Prof. Christoph Engel (Max Planck Institute for research on Collective Goods) and Prof. Cristoph Sorge (University of Saarland) for their valuable feedback during the course of this work. Additionally, Ingmar Weber is supported by funding from the Alexander von Humboldt Foundation and its founder, the Federal Ministry of Education and Research (Bundesministerium für Bildung und Forschung).
8 Ethics considerations
We placed the highest priority on ensuring that our research project adheres to standard ethical guidelines. To this end, and before initiating our study, we acquired approval from the Ethical Review Board (ERB) of our University. Before surveying a participant or obtaining their DDPs, we provided them with an online consent form and asked them to meticulously review the outlined terms in the consent form and explicitly provide us with their consent to these terms. We refrained from collecting any PII from our participants who took part in our surveys and donated us their DDPs. Specifically, we removed any available PII inside the DDPs at the front-end and also anonymize them before they are sent to our back-end server. Furthermore, according to the stated terms and conditions in the ERB approval, we took stringent measures to protect the privacy of our participants. In particular, we would not share their donated anonymized DDPs with any third parties, and we would completely remove these DDPs within a specified time after the conclusion of our study.
When conducting our surveys, we ensured that the questions did not pose any potential risks to our participants. Specifically, when surveying our participants on different representations of the DDPs, we filled out data categories based on sample DDPs that our research team obtained from their own personal accounts on the studied platforms. Moreover, we checked the contents of these data categories and verified that they do not endanger our participants in any way. We utilized bot accounts that exclusively interacted with publicly available content on the platforms to carry out one part of our experiments in Section 6. We took multiple steps to minimize the inadvertent consequences of running these bot accounts. To this extent, we intentionally distributed the incurred overhead of our experiment over multiple short browsing sessions to prevent causing any sort of disruptions to the short-format video platforms. Specifically, on a single day of our experiment, a bot account had only one video browsing session and viewed up to a maximum of 25 short-format videos. Moreover, although our bot accounts tangibly engaged with some videos by liking them, we used only three bot accounts across all the platforms, so we do not foresee any potential harm resulting from these activities. Furthermore, although the Terms of Service of the studied platforms prohibit the usage of automated bots, the benefits of our study outweigh any potential risks associated with the activities conducted by our bot accounts. Finally, we affirm that we fully conformed to ethical research standards throughout our work [28]. This includes but is not limited to, presenting our findings in aggregate form and guaranteeing the anonymity of all our participants.
9 Open science
We comply with standard open science guidelines to the fullest extent permitted by the terms and conditions outlined in the ERB approval that we obtained. To this end, we publicly share multiple materials that we utilized to carry out our research project in References. Specifically, we provide details on the surveys which we conducted in Appendix B. These details include the exact questions and choices that we gave to our participants. Moreover, we present details on the prompts we fed to the LLM and the responses that we got from the LLM also in Appendix B.
As previously stated in Section 8, according to the terms outlined in our approved ERB proposal, we would remove all the DDPs that we collected in Section 6 from our database after the completion of our research project. Therefore, we cannot make these anonymized DDPs publicly available.
Appendix A Content of DDPs provided by YouTube, TikTok, and Instagram.
Table 5 presents an overview of the content included in the DDPs offered by these three platforms. Note that although Instagram has cookies field it just shows the last four digits of the cookie name and no other details.
Information Type TikTok Instagram YouTube Minimum expected fields
User’s usage Watch Y N Y Content Id, ts
Search Y Y Y Query term, ts
Comment Y Y Y Comment text, Content Id, ts
Like Y Y $Y^{g}$ Content Id, ts
Messages Y Y NA Content, User Id, ts
Save/Favourite Y Y Y Content Id, ts
Share(In-app) Y Y NA Content Id, ts
Share(Across-app) Y N N Content Id, ts
Interests/Topics Y Y N List of topics
Time spent $N^{*}$ $N^{*}$ $N^{*}$ Duration/frequency
User’s content Media Y Y Y Media file/URL
Text details Y Y Y title
Location Y Y Y Some place identifiers
Date time Y Y Y ts
Device - Y N Device model, OS
Other user interactions $N^{*}$ $N^{*}$ $N^{*}$ Likes, Comments
Personal details Account details Y Y $Y^{g}$ Username, DOB, Email, Profile photo
Connections Y Y Y Username, ts
Login history Y Y $Y^{g}$ IP, ts
Current devices Y Y $Y^{g}$ User agent
Current camera NA Y NA Version/type
Location Y Y $Y^{g}$ Place identifiers
Account changes - Y N Type, Old, New values, ts
Advertisements Ads viewed N N Y Content Id, ts
Personlization ${N^{*}}$ N N Reasons why the ad was shown
Access to your data N Y N Which and how (in store visit etc.)
Miscellaneous Off-platform Y Y N Platform, ts, activity
Link history - Y - Link, ts
Cookies - Y - -
Table 5: Data Transparency: Overview of collected and shared information. ${N^{*}}$ denotes the details are found in the app but not in the GDPR dump. ${Y^{g}}$ denotes the details are found in google’s DDP, but not in YouTube’s DDP. (ts-TimeStamp)
Appendix B Survey questions, prompts, and responses from an LLM.
We included the following in the supplementary material [^6] https://zenodo.org/records/14721885 :
1. Survey questions used for Section 4.
2. Overall DDPs used for evaluation in Section 4.
2. Prompts provided to LLM and responses from LLM used in Section 5.
4. Survey question used for Section 5.
Appendix C Comprehensibility of DDPs
Table 6 presents the distribution of the participants recruited for the survey in Section 4.
Attribute Type Count Percentage
Gender Male 228 57.0
Female 161 40.3
Other 9 2.3
Prefer not to say 2 0.5
Age 18-25 95 23.7
25-30 104 26.0
30-40 114 28.5
40+ 87 21.8
Country France 100 25.0
Germany 100 25.0
Spain 100 25.0
Italy 100 25.0
Table 6: Distribution of participants in survey reported in Section 4 based on their self-reported gender, age, and country of residence.
C.1 Evaluation for the requirements across the platforms
Table 7 shows the data categories considered for evaluating the comprehensibility of DDPs.
Category What does it mean?
Watch history List of contents watched by a user on the platform.
Search history List of terms/users searched by a user on the platform.
Comments List of comments written by a user on others posts/videos on the platform.
Saved List of contents saved by a user for future reference on the platform.
Connections List of other users connected to a user on the platform.
Location The most recent location as recorded by the platform.
Autofill information Personally Identifiable Information including name, mail id, phone number, address etc. often collected implicitly.
Off-platform activity User’s activity data shared by other platforms with the platforms under consideration.
Devices Details of the devices that a user has used to connect to the platform.
Login history Device and network from which the user has logged into the platforms.
Like history List of contents liked by a user.
Personal information Personally Identifiable Information, including name, mail ID, phone number, address, etc., shared by the user.
User’s content List of contents uploaded or posted by the user on the platform.
Table 7: A list of categories of data that were considered in the adherence evaluation phase of the user survey.
Figure 8 shows the representation of watch history for the three platforms and sample question format.
Figure 8: Figure representing how the watch history from the three platforms was displayed to users for comparison and to evaluate the four properties: conciseness, clear and plain language, intelligibility, and transparency.
Appendix D Improved DDPs comprehensibility
Table 8 presents the distribution of the participants recruited for the survey in Section 5.
Attribute Type Count Percentage
Gender Male 100 50.0
Female 99 49.5
Prefer not to say 1 0.5
Age 18-25 54 27.0
25-30 57 28.5
30-40 57 28.5
40+ 32 16.0
Country France 50 25.0
Germany 50 25.0
Spain 50 25.0
Italy 50 25.0
Table 8: Participant distribution in the survey reported in Section 5.2 based on their self-reported gender, age, and country.